OpenStudy (anonymous):

im doing something wrong here (1^2*2)^4

OpenStudy (anonymous):

This, to me, means: \[ (1^2\cdot2)^4=(2)^4=16 \]But, I don't know if you meant \(1^{2\cdot2}\)

OpenStudy (anonymous):

how did you get the (2)^4

OpenStudy (anonymous):

By order of operations, we do (inside the parentheses) exponentiation, first, then multiply, then exponentiation, again: \[ (1^2\cdot2)^4=\left((1\cdot1\right)\cdot2)^4=(1\cdot2)^4=(2)^4=16 \]

OpenStudy (anonymous):

ok